OpenStudy (anonymous):

The density of an object is 450 kg/m3. What is the mass of the object if the volume is 2.4 m3? A. 187.5 kg B. 452.4 kg C. 1080 kg D. 6220.8 kg

jimthompson5910 (jim_thompson5910):

Use the formula D = m/v

jimthompson5910 (jim_thompson5910):

D = density m = mass v = volume

OpenStudy (anonymous):

A?

jimthompson5910 (jim_thompson5910):

D = m/v 450 = m/2.4 m = ???